OpenStudy (anonymous):

How do you solve this? (x+4)^2/3=x

OpenStudy (zehanz):

Write it as follows:\[\frac{ (x+4)^2 }{ 3 }=x\]Multiply with 3: (x+4)²=3x Then: expand the brackets, rearrange so that the right hand side = 0 Hopefully, it can be factored then.

OpenStudy (anonymous):

why is the 3 on the denominator.

OpenStudy (anonymous):

when i wrote 2/3 thats the exponent

OpenStudy (zehanz):

You should have used brackets :( Well now you have:\[(x+4)^{\frac{2}{3}}=x\] If you raise both sides to the 3rd power, you get: (x+4)²=x³. This cannot be solved easily, but fortunately there is a very simple solution. You may see it...

OpenStudy (anonymous):

haha yea i forgot sorry. well thanks